OpenStudy (anonymous):

Solve the system by substitution -4.5x-2y=-12.5 3.25x-y=-0.75 Please help! I have been stuck on this problem for a LONG time...

OpenStudy (zpupster):

you have to solve one of the equations for either x or y it looks like if we take the 2nd equation it would be easiest because y variable is all alone. 3.25x-y=-0.75 add y both sides and add .75 to both sides we get 3.25x + .75 = y now we know y, substitute that in first equation solve for x -4.5x-2(3.25x + .75)=-12.5 distribute the -2 -4.5x - 6.5x -1.5=-12.5 add 1.5 to both sides combine x terms -11x=-11 x=1 no we know x plug 1back in the equation for x-- and solve for y remember to do a check